# Bias and Fairness in Large Language Models: A Survey

### v2

@@ 5.2.2 Loss Function Modification
Adversarial learning.
In adversarial learning settings, a predictor and attacker are simultaneously trained, and the predictor aims to minimize its own loss while maximizing the attacker’s. In our setting, this training paradigm can be used to learn models that satisfy an equality constraint with respect to a protected attribute. Zhang et al. (2018a) present an early general, model-agnostic framework for bias mitigation with adversarial learning, applicable to text data. While the predictor models the desired outcome, the adversary learns to predict a protected attribute, given an equality constraint that satisfies demographic parity, equality of odds, or equality of opportunity. Other works have since followed this framework (Han et al., 2021b; Jin et al., 2021), training an encoder and discriminator, where the discriminator predicts a protected attribute from a hidden representation, and the encoder aims to prevent the discriminator from discerning these protected attributes from the encodings.
Several works have proposed improvements to this general framework. For bias mitigation in a setting with only limited labeling of protected attributes, Han et al. (2021a) propose a modified optimization objective that separates discriminator training from the main model training, so that the discriminator can be selectively applied to only the instances with a social group label. To model equal opportunity (as opposed to demographic parity, where the target label and social group are independent), Han et al. (2022b) add an augmentation layer between the encoder and predicted attribute classifier, allowing the discriminator to access the target label. Rekabsaz et al. (2021) adapt these methods to the ranking of information retrieval results to reduce bias while maintaining relevance, proposing a gender-invariant ranking model called AdvBERT. Contrastive pairs consist of a relevant and non-relevant document to a query, with a corresponding social group label denoting if the query or document contains the protected attribute. The adversarial discriminator predicts the social group label from an encoder, while the encoder simultaneously tries to trick the discriminator while also maximizing relevance scores.
Adversarial learning can also be used to adversarially attack a model during training. Wang et al. (2021) propose to remove bias information from pre-trained embeddings for some downstream classification task by generating adversarial examples with a protected attribute classifier. The authors generate worst-case representations by perturbing the embeddings to maximize the loss of the protected attribute classifier, which are used for training.
Reinforcement learning.
@@ 5.2.3 Selective Parameter Updating
Gira et al. (2022) freeze over 99% of a model’s parameters before fine-tuning on the WinoBias (Zhao et al., 2019) and CrowS-Pairs (Nangia et al., 2020) datasets, only updating a selective set of parameters, such as layer norm parameters or word positioning embeddings. Ranaldi et al. (2023) only update the attention matrices of the pre-trained model and freeze all other parameters for fine-tuning on the PANDA (Qian et al., 2022) dataset. Instead of unfreezing a pre-determined set of parameters, Yu et al. (2023a) only optimize weights with the greatest contributions to bias within a domain, with gender-profession demonstrated as an example. Model weights are rank-ordered and selected based on the gradients of contrastive sentence pairs differing along some demographic axis.
@@ 5.2.4 Filtering Model Parameters
Besides fine-tuning techniques that simply update model parameters to reduce bias, there are also techniques focused on filtering or removing specific parameters (e.g., by setting them to zero) either during or after the training or fine-tuning of the model. Joniak & Aizawa (2022) use movement pruning, a technique that removes some weights of a neural network, to select a least-biased subset of weights from the attention heads of a pre-trained model. During fine-tuning, they freeze the weights and independently optimize scores with a debiasing objective. The scores are thresholded to determine which weights to remove.
@@ 5.2.5 Discussion and Limitations
Beyond computational limitations, in-training mitigations target different modeling mechanisms, which may vary their effectiveness. For instance, given the weak relationship between biases in the embedding space and biases in downstream tasks as discussed in Section 3.3.3, embedding-based loss function modifications may have limited effectiveness. On the other hand, since attention may be one of the primary ways that bias is encoded in LLMs (Jeoung & Diesner, 2022), attention-based loss function modifications may be more effective. Future research can better understand which components of LLMs encode, reproduce, and amplify bias to enable more targeted in-training mitigations.
Finally, the form of the loss function, or the reward given in reinforcement learning, implicitly assumes some definition of fairness, most commonly some notion of invariance with respect to social groups. These assumptions should be stated explicitly. Moreover, future work can propose alternative loss functions to capture a broader scope of fairness desiderata, which should be tailored to specific downstream applications and settings.
@@ 5.3 Intra-Processing Mitigation
Following Savani et al. (2020)’s definition, we consider intra-processing methods to be those that take a pre-trained, perhaps fine-tuned, model as input, and modify the model’s behavior without further training or fine-tuning to generate debiased predictions at inference; as such, these techniques may also be considered to be inference stage mitigations. Intra-processing techniques include decoding strategies that change the output generation procedure, post hoc model parameter modifications, and separate debiasing networks that can be applied modularly during inference. Examples are shown in Figure 9.
Figure 9: Example Intra-Processing Mitigation Techniques (§ 5.3). We show several methods that modify a model’s behavior without training or fine-tuning. Constrained next-token search may prohibit certain outputs during beam search (e.g., a derogatory term "@&!", in this example), or generate and rerank alternative outputs (e.g., "he" replaced with "she"). Modified token distribution redistributes next-word probabilities to produce more diverse outputs and avoid biased tokens. Weight distribution, in this example, illustrates how post hoc modifications to attention matrices may narrow focus to less stereotypical tokens (Zayed et al., 2023b). Modular debiasing networks fuse the main LLM with stand-alone networks that can remove specific dimensions of bias, such as gender or racial bias.
@@ 5.3.1 Decoding Strategy Modification
Constrained next-token search.
Constrained next-token search considers methods that change the ranking of the next token by adding additional requirements. In a simple and coarse approach, Gehman et al. (2020) and Xu et al. (2020) propose word- or $n$-gram blocking during decoding, prohibiting the use of tokens from an unsafe word list. However, biased outputs can still be generated from a set of unbiased tokens or $n$-grams. To improve upon token blocking strategies, more nuanced approaches constrain text generation by comparing the most likely or a potentially-biased generation to a counterfactual or less biased version. Using a counterfactual-based method, Saunders et al. (2022) use a constrained beam search to generate more gender-diverse outputs at inference. The constrained beam search generates an $n$-best list of outputs in two passes, first generating the highest likelihood output and then searching for differently-gendered versions of the initial output. Comparing instead to known biases in the data, Sheng et al. (2021a) compare $n$-gram features from the generated outputs with frequently-occurring biased (or otherwise negative) demographically-associated phrases in the data. These $n$-gram features constrain the next token prediction by requiring semantic similarity with unbiased phrases and dissimilarity with biased phrases. Meade et al. (2023) compare generated outputs to safe example responses from similar contexts, reranking candidate responses based on their similarity to the safe example. Instead of comparing various outputs, Lu et al. (2021) more directly enforce lexical constraints given by predicate logic statements, which can require the inclusion or exclusion of certain tokens. The logical formula is integrated as a soft penalty during beam search.
@@
Modified token distribution.
Changing the distribution from which tokens are sampled can increase the diversity of the generated output or enable the sampling of less biased outputs with greater probability. Chung et al. (2023) propose two decoding strategies to increase diversity of generated tokens. Logit suppression decreases the probability of generating already-used tokens from previous generations, which encourages the selection of lower-frequency tokens. Temperature sampling flattens the next-word probability distribution to also encourage the selection of less-likely tokens. Kim et al. (2023) also modify the output token distribution using reward values obtained from a toxicity evaluation model. The authors raise the likelihood of tokens that increase a reward value, and lower ones that do not. Gehman et al. (2020) similarly increases the likelihood of non-toxic tokens, adding a (non-)toxicity score to the logits over the vocabulary before normalization. Liu et al. (2023b) alternatively redistribute the probability mass with bias terms. The proposed method seeks to minimize a constraint function such as toxicity with an iterative sequence generation process, tuning bias terms added to the predicted logits at each decoding step. After decoding for several steps, the bias terms are updated with gradient descent to minimize the toxicity of the generated sequence.
Another class of approaches modifies token probabilities by comparing two outputs differing in their level of bias. Liu et al. (2021a) uses a combination of a pre-trained model and two smaller language models during decoding, one expert that models non-toxic text, and one anti-expert that models toxic text. The pre-trained logits are modified to increase the probability of tokens with high probability under the expert and low probability under the anti-expert. Hallinan et al. (2023) similarly identify potentially toxic tokens with an expert and an anti-expert, and mask and replace candidate tokens with less toxic alternatives. In GeDi, Krause et al. (2021) also compares the generated outputs from two language models, one conditioned on an undesirable attribute like toxicity, which guides each generation step to avoid toxic words. Instead of using an additional model, Schick et al. (2021) propose a self-debiasing framework. The authors observe that pre-trained models can often recognize their own biases in the outputs they produce and can describe these behaviors in their own generated descriptions. This work compares the distribution of the next word given the original input, to the distribution given the model’s own reasoning about why the input may be biased. The model chooses words with a higher probability of being unbiased.
@@ 5.3.2 Weight Redistribution
The weights of a trained model may be modified post hoc without further training. Given the potential associations between attention weights and encoded bias (Jeoung & Diesner, 2022), redistributing attention weights may change how the model attends to biased words or phrases. Though Attanasio et al. (2022) and Gaci et al. (2022) propose in-training approaches (see Section 5.2.2), Zayed et al. (2023a) modify the attention weights after training, applying temperature scaling controlled by a hyperparameter that can be tuned to maximize some fairness metric. The hyperparameter can either increase entropy to focus on a broader set of potentially less stereotypical tokens, or can decrease entropy to attend to a narrower context, which may reduce exposure to stereotypical tokens.
@@ 5.3.3 Modular Debiasing Networks
Hauzenberger et al. (2023) propose a technique that trains several subnetworks that can be applied modularly at inference time to remove a specific set of biases. The work adapts diff pruning (Guo et al., 2021) to the debiasing setting, mimicking the training of several parallel models debiased along different dimensions, and storing changes to the pre-trained model’s parameters in sparse subnetworks. The output of this technique is several stand-alone modules, each corresponding to a debiasing task, that can be used with a base pre-trained model during inference. Similarly, Kumar et al. (2023a) introduce adapter modules for bias mitigation, based on adapter networks that learn task-specific parameters (Pfeiffer et al., 2021). This work creates an adapter network by training a single-layer multilayer perceptron with the objective of removing protected attributes, with an additional fusion module to combine the original pre-trained model with the adapter.
@@ 5.3.4 Discussion and Limitations
The primary limitations of intra-processing mitigations center on decoding strategy modifications; work in weight redistribution and modular debiasing networks for bias mitigation is limited, and future work can expand research in these areas. One of the biggest challenges in decoding strategy modifications is balancing bias mitigation with diverse output generation. These methods typically rely on identifying toxic or harmful tokens, which requires a classification method that is not only accurate but also unbiased in its own right (see Section 3.5.4 for discussion of challenges with classifier-based techniques). Unfortunately, minority voices are often disproportionately filtered out as a result. For instance, Xu et al. (2021) find that techniques that reduce toxicity can in turn amplify bias by not generating minority dialects like African-American English. Any decoding algorithm that leverages some heuristic to identify bias must take special care to not further marginalize underrepresented and minoritized voices. Kumar et al. (2023b) also warn that decoding algorithms may be manipulated to generate biased language.
@@ 5.4 Post-Processing Mitigation
Figure 10: Example Post-processing Mitigation Techniques (§ 5.4). We illustrate how post-processing methods can replace a gendered output to a gender-neutral version. Keyword replacement methods first identify protected attribute terms (i.e., "mothers", "he"), and then generate an alternative output. Machine translation methods train a neural machine translator on a parallel biased-unbiased corpus and feed the original output into the model to produce an unbiased output.
@@ 5.4.1 Rewriting
Keyword replacement.
Keyword replacement approaches aim to identify biased tokens and predict replacements, while preserving the content and style of the original output. Tokpo & Calders (2022) use LIME (Ribeiro et al., 2016) to identify tokens responsible for bias in an output and predict new tokens for replacement based on the latent representations of the original sentence. Dhingra et al. (2023) employ SHAP (Lundberg & Lee, 2017) to identify stereotypical words towards queer people, providing reasoning for why the original word was harmful. They then re-prompt the language model to replace those words, using style transfer to preserve the semantic meaning of the original sentence. He et al. (2021) detect and mask protected attribute tokens using a protected attribute classifier, and then apply a neural rewriting model that takes in the masked sentence as input and regenerates the output without the protected attribute.
Machine translation.
Another class of rewriter models translates from a biased source sentence to a neutralized or un-based target sentence. This can be framed as a machine translation task, training on parallel corpora that translates from a biased (e.g., gendered) to an unbiased (e.g., gender-neutral or opposite gender) alternative. To provide gender-neutral alternatives to sentences with gendered pronouns, several works (Jain et al., 2021; Sun et al., 2021; Vanmassenhove et al., 2021) use a rules-based approach to generate parallel debiased sentences from biased sources, and then train a machine translation model to translate from biased sentences to debiased ones. Instead of generating a parallel corpus using biased sentences as the source, Amrhein et al. (2023) leverage backward augmentation to filter through large corpora for gender-fair sentences, and then add bias to generate artificial source sentences.
@@
Other neural rewriters.
Ma et al. (2020) focus specifically on editing the power dynamics and agency levels encoded in verbs, proposing a neural model that can reconstruct and paraphrase its input, while boosting the use of power- or agency-connoted words. Majumder et al. (2022) present InterFair for user-informed output modification during inference. After scoring words important for task prediction and words associated with bias, the user can critique an adjust the scores to inform rewriting.

### v3

@@ 5.2.2 Loss Function Modification
Adversarial learning
In adversarial learning settings, a predictor and attacker are simultaneously trained, and the predictor aims to minimize its own loss while maximizing the attacker’s. In our setting, this training paradigm can be used to learn models that satisfy an equality constraint with respect to a protected attribute. Zhang, Lemoine, and Mitchell (2018) present an early general, model-agnostic framework for bias mitigation with adversarial learning, applicable to text data. While the predictor models the desired outcome, the adversary learns to predict a protected attribute, given an equality constraint (e.g., demographic parity, equality of odds, or equal opportunity). Other works have since followed this framework (Han, Baldwin, and Cohn, 2021b; Jin et al., 2021), training an encoder and discriminator, where the discriminator predicts a protected attribute from a hidden representation, and the encoder aims to prevent the discriminator from discerning these protected attributes from the encodings.
Several works have proposed improvements to this general framework. For bias mitigation in a setting with only limited labeling of protected attributes, Han, Baldwin, and Cohn (2021a) propose a modified optimization objective that separates discriminator training from the main model training, so that the discriminator can be selectively applied to only the instances with a social group label. For more complete dependence between the social group and outcome, Han, Baldwin, and Cohn (2022b) add an augmentation layer between the encoder and predicted attribute classifier and allow the discriminator to access the target label. Rekabsaz, Kopeinik, and Schedl (2021) adapt these methods to the ranking of information retrieval results to reduce bias while maintaining relevance, proposing a gender-invariant ranking model called AdvBERT. Contrastive pairs consist of a relevant and non-relevant document to a query, with a corresponding social group label denoting if the query or document contains the protected attribute. The adversarial discriminator predicts the social group label from an encoder, while the encoder simultaneously tries to trick the discriminator while also maximizing relevance scores.
Adversarial learning can also be used to adversarially attack a model during training. Wang et al. (2021) propose to remove bias information from pre-trained embeddings for some downstream classification task by generating adversarial examples with a protected attribute classifier. The authors generate worst-case representations by perturbing and training on embeddings that maximize the loss of the protected attribute classifier.
Reinforcement learning
@@ 5.2.3 Selective Parameter Updating
Gira, Zhang, and Lee (2022) freeze over 99% of a model’s parameters before fine-tuning on the WinoBias (Zhao et al., 2019) and CrowS-Pairs (Nangia et al., 2020) datasets, only updating a selective set of parameters, such as layer norm parameters or word positioning embeddings. Ranaldi et al. (2023) only update the attention matrices of the pre-trained model and freeze all other parameters for fine-tuning on the PANDA (Qian et al., 2022) dataset. Instead of unfreezing a pre-determined set of parameters, Yu et al. (2023a) only optimize weights with the greatest contributions to bias within a domain, with gender-profession demonstrated as an example. Model weights are rank-ordered and selected based on the gradients of contrastive sentence pairs differing along some demographic axis.
@@ 5.2.4 Filtering Model Parameters
Besides fine-tuning techniques that simply update model parameters to reduce bias, there are also techniques focused on filtering or removing specific parameters (e.g., by setting them to zero) either during or after the training or fine-tuning of the model. Joniak and Aizawa (2022) use movement pruning (Sanh, Wolf, and Rush, 2020), a technique that removes some weights of a neural network, to select a least-biased subset of weights from the attention heads of a pre-trained model. During fine-tuning, they freeze the weights and independently optimize scores with a debiasing objective. The scores are thresholded to determine which weights to remove. To build robustness against the circumvention of safety alignment ("jailbreaking"), including resistance to hate speech and discriminatory generations, Hasan, Rugina, and Wang (2024) alternatively use WANDA (Sun et al., 2023b), which induces sparsity by pruning weights with a small element-wise product between the weight matrix and input feature activations, as a proxy for low-importance parameters. The authors show that pruning 10-20% of model parameters increases resistance to jailbreaking, but more extensive pruning can have detrimental effects.
Proskurina, Metzler, and Velcin (2023) provide further evidence that aggressive pruning can have adverse effects: for hate speech classification, models with pruning of 30% or more of the original parameters demonstrate increased levels of gender, race, and religious bias. In an analysis of stereotyping and toxicity classification in text, Ramesh et al. (2023) also find that pruning may amplify bias in some cases, but with mixed effects and dependency on the degree of pruning.
@@ 5.2.5 Discussion and Limitations
Beyond computational limitations, in-training mitigations target different modeling mechanisms, which may vary their effectiveness. For instance, given the weak relationship between biases in the embedding space and biases in downstream tasks as discussed in Section 3.3.3, embedding-based loss function modifications may have limited effectiveness. On the other hand, since attention may be one of the primary ways that bias is encoded in LLMs (Jeoung and Diesner, 2022), attention-based loss function modifications may be more effective. Future research can better understand which components of LLMs encode, reproduce, and amplify bias to enable more targeted in-training mitigations.
Finally, the form of the loss function, or the reward given in reinforcement learning, implicitly assumes some definition of fairness, most commonly some notion of invariance with respect to social groups, even though harms often operate in nuanced and distinct ways for various social groups. Treating social groups or their outcomes as interchangeable ignores the underlying forces of injustice. The assumptions encoded in the choice of loss function should be stated explicitly. Moreover, future work can propose alternative loss functions to capture a broader scope of fairness desiderata, which should be tailored to specific downstream applications and settings.
@@ 5.3 Intra-Processing Mitigation
Following Savani, White, and Govindarajulu (2020)’s definition, we consider intra-processing methods to be those that take a pre-trained, perhaps fine-tuned, model as input, and modify the model’s behavior without further training or fine-tuning to generate debiased predictions at inference; as such, these techniques may also be considered to be inference stage mitigations. Intra-processing techniques include decoding strategies that change the output generation procedure, post hoc model parameter modifications, and separate debiasing networks that can be applied modularly during inference. Examples are shown in Figure 9.
Figure 9: Example Intra-Processing Mitigation Techniques (§ 5.3). We show several methods that modify a model’s behavior without training or fine-tuning. Constrained next-token search may prohibit certain outputs during beam search (e.g., a derogatory term "@&!," in this example), or generate and rerank alternative outputs (e.g., "he" replaced with "she"). Modified token distribution redistributes next-word probabilities to produce more diverse outputs and avoid biased tokens. Weight distribution, in this example, illustrates how post hoc modifications to attention matrices may narrow focus to less stereotypical tokens (Zayed et al., 2023b). Modular debiasing networks fuse the main LLM with stand-alone networks that can remove specific dimensions of bias, such as gender or racial bias.
@@ 5.3.1 Decoding Strategy Modification
Constrained next-token search
Constrained next-token search considers methods that change the ranking of the next token by adding additional requirements. In a simple and coarse approach, Gehman et al. (2020) and Xu et al. (2020) propose word- or $n$-gram blocking during decoding, prohibiting the use of tokens from an offensive word list. However, biased outputs can still be generated from a set of unbiased tokens or $n$-grams. To improve upon token-blocking strategies, more nuanced approaches constrain text generation by comparing the most likely or a potentially-biased generation to a counterfactual or less biased version. Using a counterfactual-based method, Saunders, Sallis, and Byrne (2022) use a constrained beam search to generate more gender-diverse outputs at inference. The constrained beam search generates an $n$-best list of outputs in two passes, first generating the highest likelihood output and then searching for differently-gendered versions of the initial output. Comparing instead to known biases in the data, Sheng et al. (2021a) compare $n$-gram features from the generated outputs with frequently-occurring biased (or otherwise negative) demographically-associated phrases in the data. These $n$-gram features constrain the next token prediction by requiring semantic similarity with unbiased phrases and dissimilarity with biased phrases. Meade et al. (2023) compare generated outputs to safe example responses from similar contexts, reranking candidate responses based on their similarity to the safe example. Instead of comparing various outputs, Lu et al. (2021) more directly enforce lexical constraints given by predicate logic statements, which can require the inclusion or exclusion of certain tokens. The logical formula is integrated as a soft penalty during beam search.
@@
Modified token distribution
Changing the distribution from which tokens are sampled can increase the diversity of the generated output or enable the sampling of less biased outputs with greater probability. Chung, Kamar, and Amershi (2023) propose two decoding strategies to increase diversity of generated tokens. Logit suppression decreases the probability of generating already-used tokens from previous generations, which encourages the selection of lower-frequency tokens. Temperature sampling flattens the next-word probability distribution to also encourage the selection of less-likely tokens. Kim et al. (2023) also modify the output token distribution using reward values obtained from a toxicity evaluation model. The authors raise the likelihood of tokens that increase a reward value, and lower ones that do not. Gehman et al. (2020) similarly increases the likelihood of non-toxic tokens, adding a (non-)toxicity score to the logits over the vocabulary before normalization. Liu, Khalifa, and Wang (2023) alternatively redistribute the probability mass with bias terms. The proposed method seeks to minimize a constraint function such as toxicity with an iterative sequence generation process, tuning bias terms added to the predicted logits at each decoding step. After decoding for several steps, the bias terms are updated with gradient descent to minimize the toxicity of the generated sequence.
Another class of approaches modifies token probabilities by comparing two outputs differing in their level of bias. Liu et al. (2021a) uses a combination of a pre-trained model and two smaller language models during decoding, one expert that models non-toxic text, and one anti-expert that models toxic text. The pre-trained logits are modified to increase the probability of tokens with high probability under the expert and low probability under the anti-expert. Hallinan et al. (2023) similarly identify potentially toxic tokens with an expert and an anti-expert, and mask and replace candidate tokens with less toxic alternatives. In GeDi, Krause et al. (2021) also compares the generated outputs from two language models, one conditioned on an undesirable attribute like toxicity, which guides each generation step to avoid toxic words. Instead of using an additional model, Schick, Udupa, and Schütze (2021) propose a self-debiasing framework. The authors observe that pre-trained models can often recognize their own biases in the outputs they produce and can describe these behaviors in their own generated descriptions. This work compares the distribution of the next word given the original input, to the distribution given the model’s own reasoning about why the input may be biased. The model chooses words with a higher probability of being unbiased.
@@ 5.3.2 Weight Redistribution
The weights of a trained model may be modified post hoc without further training. Given the potential associations between attention weights and encoded bias (Jeoung and Diesner, 2022), redistributing attention weights may change how the model attends to biased words or phrases. Though Attanasio et al. (2022) and Gaci et al. (2022) propose in-training approaches (see Section 5.2.2), Zayed et al. (2023a) modify the attention weights after training, applying temperature scaling controlled by a hyperparameter that can be tuned to maximize some fairness metric. The hyperparameter can either increase entropy to focus on a broader set of potentially less stereotypical tokens, or can decrease entropy to attend to a narrower context, which may reduce exposure to stereotypical tokens.
@@ 5.3.3 Modular Debiasing Networks
Hauzenberger et al. (2023) propose a technique that trains several subnetworks that can be applied modularly at inference time to remove a specific set of biases. The work adapts diff pruning (Guo, Rush, and Kim, 2021) to the debiasing setting, mimicking the training of several parallel models debiased along different dimensions, and storing changes to the pre-trained model’s parameters in sparse subnetworks. The output of this technique is several stand-alone modules, each corresponding to a debiasing task, that can be used with a base pre-trained model during inference. Similarly, Kumar et al. (2023a) introduce adapter modules for bias mitigation, based on adapter networks that learn task-specific parameters (Pfeiffer et al., 2021). This work creates an adapter network by training a single-layer multilayer perceptron with the objective of removing protected attributes, with an additional fusion module to combine the original pre-trained model with the adapter.
@@ 5.3.4 Discussion and Limitations
The primary limitations of intra-processing mitigations center on decoding strategy modifications; work in weight redistribution and modular debiasing networks for bias mitigation is limited, and future work can expand research in these areas. One of the biggest challenges in decoding strategy modifications is balancing bias mitigation with diverse output generation. These methods typically rely on identifying toxic or harmful tokens, which requires a classification method that is not only accurate but also unbiased in its own right (see Section 3.5.4 for discussion of challenges with classifier-based techniques). Unfortunately, minority voices are often disproportionately filtered out as a result. For instance, Xu et al. (2021) find that techniques that reduce toxicity can in turn amplify bias by not generating minority dialects like African-American English. Any decoding algorithm that leverages some heuristic to identify bias must take special care to not further marginalize underrepresented and minoritized voices. Kumar et al. (2023b) also warn that decoding algorithms may be manipulated to generate biased language by increasing, rather than decreasing, the generation of toxic or hateful text.
@@ 5.4 Post-Processing Mitigation
Figure 10: Example Post-processing Mitigation Techniques (§ 5.4). We illustrate how post-processing methods can replace a gendered output with a gender-neutral version. Keyword replacement methods first identify protected attribute terms (i.e., "mothers," "he"), and then generate an alternative output. Machine translation methods train a neural machine translator on a parallel biased-unbiased corpus and feed the original output into the model to produce an unbiased output.
@@ 5.4.1 Rewriting
Keyword replacement
Keyword replacement approaches aim to identify biased tokens and predict replacements, while preserving the content and style of the original output. Tokpo and Calders (2022) use LIME (Ribeiro, Singh, and Guestrin, 2016) to identify tokens responsible for bias in an output and predict new tokens for replacement based on the latent representations of the original sentence. Dhingra et al. (2023) employ SHAP (Lundberg and Lee, 2017) to identify stereotypical words towards queer people, providing reasoning for why the original word was harmful. They then re-prompt the language model to replace those words, using style transfer to preserve the semantic meaning of the original sentence. He, Majumder, and McAuley (2021) detect and mask protected attribute tokens using a protected attribute classifier, and then apply a neural rewriting model that takes in the masked sentence as input and regenerates the output without the protected attribute.
Machine translation
Another class of rewriter models translates from a biased source sentence to a neutralized or un-based target sentence. This can be framed as a machine translation task, training on parallel corpora that translates from a biased (e.g., gendered) to an unbiased (e.g., gender-neutral or opposite gender) alternative. To provide gender-neutral alternatives to sentences with gendered pronouns, several works (Jain et al., 2021; Sun et al., 2021; Vanmassenhove, Emmery, and Shterionov, 2021) use a rules-based approach to generate parallel debiased sentences from biased sources, and then train a machine translation model to translate from biased sentences to debiased ones. Instead of generating a parallel corpus using biased sentences as the source, Amrhein et al. (2023) leverage backward augmentation to filter through large corpora for gender-fair sentences, and then add bias to generate artificial source sentences.
@@
Other neural rewriters
Ma et al. (2020) focus specifically on editing the power dynamics and agency levels encoded in verbs, proposing a neural model that can reconstruct and paraphrase its input, while boosting the use of power- or agency-connoted words. Majumder, He, and McAuley (2022) present InterFair for user-informed output modification during inference. After scoring words important for task prediction and words associated with bias, the user can critique and adjust the scores to inform rewriting.